# Improving Automated Feedback Systems for Tutor Training in Low-Resource Scenarios through Data Augmentation

### v1

@@ IV-A2 Analysis of Results
The average M-IoU, IoU, $F_{2}$ score performance of the fine-tuned models, trained with the training set split with five different random seeds, is depicted in Table III and Table IV.
TABLE III: Average Scores and Standard Errors for M-IoU, IoU, and $F_{2}$ Score (effort-based praise). 520* denotes traditional augmentation methods (e.g., synonym replacement, random word insertion) with a set size of 520.
@@
13 $0.506$ $0.024$ $0.443$ $0.011$ $0.528$ $0.024$
26 $0.501$ $0.020$ $0.440$ $0.009$ $0.525$ $0.019$
65 $0.495$ $0.021$ $0.436$ $0.009$ $0.512$ $0.021$
130 $0.551$ $0.018$ $0.488$ $0.013$ $0.573$ $0.020$
260 $0.596$ $0.008$ $0.551$ $0.004$ $0.619$ $0.008$
520 $0.601$ $0.006$ $0.554$ $0.007$ $0.622$ $0.005$
520* $0.456$ $0.036$ $0.392$ $0.014$ $0.477$ $0.036$
TABLE IV: Average Scores and Standard Errors for M-IoU, IoU, and $F_{2}$ Score (outcome-based praise). 520* denotes traditional augmentation methods (e.g., synonym replacement, random word insertion) with a set size of 520.
@@
Table III presents the scores and errors for effort-based praise across different augmented training set sizes, ranging from 13 to 520 labeled responses, while Table IV details the corresponding metrics for outcome-based praise. In both cases, as the training set size increases, there is a noticeable improvement in all three metrics—M-IoU, IoU, and $F_{2}$ scores. For effort-based praise, the M-IoU score starts at 0.506 with a set size of 13 and gradually increases to 0.601 with a set size of 520, with similar trends observed for the IoU and $F_{2}$ scores, indicating that larger augmented training sets contribute to better model performance. Similarly, for outcome-based praise, the M-IoU score begins at 0.623 for the smallest set size of 13 and reaches 0.772 with the largest set size of 520, with consistent improvements also seen in the IoU and $F_{2}$ scores. These results reinforce the effectiveness of data augmentation for enhancing the model’s ability to accurately identify both effort-based and outcome-based praise. In addition to our primary methods, we employed traditional NLP data augmentation methods—random insertion, swap, deletion, and synonym replacement—to expand the original 13 responses into a dataset of 520 entries. Specifically, each original response was augmented by applying these traditional methods collectively rather than generating separate datasets for each method individually. For example, a single tutor response would undergo synonym replacement, random word insertion, swapping, and deletion in combination, producing several augmented responses simultaneously. This combined augmentation approach allowed for more extensive dataset expansion by creating more diverse variants of each original response. The augmented dataset, containing a total of 520 responses, is labeled as ’520*’ in Table III and Table IV, where we report the performance of models trained on this expanded dataset. However, these traditional data augmentation methods did not enhance performance and, in fact, resulted in worse outcomes compared to using no augmentation at all. This lack of improvement could be attributed to the introduction of incoherent or invalid sentences by these methods. For instance, the sentence “Hey Kevin, you did a great job.” was altered to “Hey, Kevin, you the did an good work,” which illustrates the potential for reduced clarity and grammatical correctness.
@@
Since expert annotations serve as the ground truth, it is not possible to directly calculate an M-IoU score for expert-level performance. However, we estimated this score using a simple method based on our previous work [6]. In that study, we validated M-IoU as an evaluation metric by engaging two trained human raters to assess the correctness of ChatGPT’s annotations, using a scale from 1 (completely incorrect) to 5 (completely correct). We calculated the correlation between these human ratings and the automatically computed M-IoU scores, observing a strong positive correlation. This result demonstrated that M-IoU aligns well with human judgment regarding annotation quality.
@@ IV-B1 Experimental Setup and Data Imbalance
As discussed in Section III-A, we have a limited dataset with only 10 labeled responses featuring person-based praise. We combine these with our existing 129 labeled responses that do not include person-based praise, which results in a total of 139 labeled responses. Due to the significant data imbalance, where the majority of the praise is effort-based or outcome-based and only a small portion is person-based, we approach this as a two-class NER problem—instead of distinguishing between effort-based praise and outcome-based praise as in RQ1, we only require the model to identify person-based praise (labeled as P) and other parts (labeled as O). If we were to fine-tune the model on all types of praise without addressing this imbalance, the model might be biased towards predicting effort-based or outcome-based praise, given their predominance in the dataset. By framing it as a two-class NER problem, we can focus on distinguishing between responses that contain person-based praise and those that do not, helping to mitigate the model’s tendency to favor the more common categories. Following the method used in RQ1, we split the dataset into a training set and a test set. The training set consists of 70 labeled responses, including 5 person-based praise responses, while the test set contains 69 labeled responses, also including 5 person-based praise responses. We then apply our data augmentation method to the training set. The augmentation multipliers were set to ×1,×2,×3$\times$ 1$\times$ 2$\times$ 3× 1 , × 2 , × 3, and ×$\times 5$ based on the model performance.
@@ IV-B2 Generalization Performance
The results of this experiment are presented in Table V.
TABLE V: Average Scores and Standard Errors for M-IoU, IoU, and $F_{2}$ Score (person-based praise). Aug. Size represents Augmented Training Set Size, and Mult. represents Augmentation Multiplier.
@@
×$\times 1$ 70 $0.808$ $0.014$ $0.797$ $0.016$ $0.810$ $0.014$
×$\times 2$ 140 $0.897$ $0.013$ $0.895$ $0.014$ $0.902$ $0.014$
×$\times 3$ 210 $0.966$ $0.017$ $0.965$ $0.013$ $0.974$ $0.018$
×$\times 5$ 350 $0.969$ $0.016$ $0.969$ $0.017$ $0.976$ $0.016$
As is shown in Table V, as the augmented training dataset size increases from 70 to 350, there is a clear upward trend across all three performance metrics: M-IoU, IoU, and $F_{2}$ Score. Specifically, the M-IoU score increases from 0.808 at a size of 70 to 0.969 at a size of 350. Similarly, the IoU score improves from 0.797 to 0.969, and the $F_{2}$ Score rises from 0.810 to 0.976. These results indicate that as we augment the training dataset, the model’s ability to identify person-based praise improves consistently across all measured metrics. This gradual improvement demonstrates the effectiveness of our data augmentation method in enhancing the model’s performance.
Specifically, when the data augmentation multiplier is set to ×$\times 3$ (namely when the augmented training dataset size is 210), the model achieves satisfactory performance, the M-IoU, IoU, and $F_{2}$ scores already reached 0.965 or higher, and even achieved a perfect score of 1 under certain random seeds, suggesting that the model was nearing its performance ceiling, which demonstrates the model’s robust ability to generalize to person-based praise. Beyond this multiplier, at ×$\times 5$, the performance plateaus, where increasing the multiplier to ×$\times 5$ yielded only marginal improvements, suggesting that further augmentation does not yield additional benefits.
The relatively high performance achieved with smaller augmentation multipliers (e.g., ×$\times 3$, namely the augmented training dataset size is 210) can be attributed to the nature of person-based praise, which often consists of straightforward adjectives describing a person’s characteristics (e.g., “You are smart.”, “You are the smartest student.”). The simplicity in the vocabulary and word types—predominantly adjectives—makes it easier for the model to recognize and tag person-based praise, in contrast to the more nuanced effort-based and outcome-based praise categories examined in RQ1. This straightforward linguistic structure allows the model to detect patterns with greater ease, as opposed to the complexity found in the other categories, where varied and context-dependent language may pose more challenges for accurate tagging.
@@ V Ablation Study
To assess whether more advanced LLMs, such as GPT-4o, provide better data augmentation performance compared to GPT-3.5, we first evaluate the performance of data augmentation using GPT-4o and GPT-3.5, keeping all other settings consistent with the previous sections. The results are presented in Table VI.
As shown in Table VI, augmenting the dataset with GPT-4o consistently outperforms augmentation with GPT-3.5 across all metrics (M-IoU, IoU, and $F_{2}$ Score). This improvement can be explained by GPT-4o’s ability to generate more valid and contextually appropriate synonym replacements compared to GPT-3.5. Additionally, we observed that GPT-3.5 often fails to adhere strictly to the clean answer structure we set in the prompt, leading to inconsistencies in the output format and introducing some noise into the augmented data. This issue was significantly less prevalent in outputs generated by GPT-4o, which produced more reliable and coherent augmentations. For example, when augmenting a sentence like “You did a good job,” GPT-3.5 might generate expressions such as “Sure, here is the synonym: You executed a commendable operation!” which may introduce the noise “Sure, here is the synonym:”.
TABLE VI: Average M-IoU, IoU, and $F_{2}$ Score of augmentation with GPT-4o and GPT-3.5. We keep the base model as GPT-3.5.
M-IoU IoU $F_{2}$ Score
Set Size GPT-4o GPT-3.5 GPT-4o GPT-3.5 GPT-4o GPT-3.5
13 $0.565$ $0.554$ $0.520$ $0.493$ $0.586$ $0.575$
26 $0.576$ $0.569$ $0.515$ $0.508$ $0.595$ $0.583$
65 $0.584$ $0.574$ $0.539$ $0.521$ $0.601$ $0.589$
130 $0.631$ $0.598$ $0.579$ $0.546$ $0.647$ $0.612$
260 $0.682$ $0.601$ $0.652$ $0.550$ $0.697$ $0.615$
520 $0.687$ $0.609$ $0.654$ $0.561$ $0.700$ $0.621$
@@ V-B Initial Training Dataset Size
In this section, we investigate the impact of different initial training dataset sizes on model performance after data augmentation. We use GPT-4o for data augmentation while keeping the base model as GPT-3.5. The results are presented in Table VII, which shows the average M-IoU for training sets with different initial sizes when augmented to size 130, 260, and 520.
TABLE VII: Average M-IoU for Training Sets with Different Initial Sizes after Augmentation.
@@
As shown in Table VII, increasing the initial training set size leads to improved performance after augmentation, but the gains diminish as the augmented dataset size grows larger. For instance, starting with an initial set size of 13, the M-IoU improves from 0.631 to 0.687 as the augmented set size increases from 130 to 520. However, when starting with an initial set size of 65, the performance gains become minimal, suggesting that when the initial set size is relatively large, further augmentation may not be necessary or only a little augmentation is required. Beyond a certain point, additional augmentation yields diminishing returns. This observation indicates that while data augmentation is effective in low-resource scenarios (e.g., starting with only 13 or 26 labeled instances), its impact becomes less pronounced as more labeled data is available initially. Therefore, for small datasets, augmentation plays a crucial role in boosting model performance, but its utility decreases as the dataset grows larger.
@@ V-C Base Model
In this section, we compare the performance of GPT-3.5 and GPT-4o as base models, while keeping GPT-4o as the augmentation model. The results are shown in Table VIII.
TABLE VIII: Average M-IoU, IoU, and $F_{2}$ Score with GPT-4o and GPT-3.5 as base model. We keep augmentation model as GPT-4o.
M-IoU IoU $F_{2}$ Score
Set Size GPT-3.5 GPT-4o GPT-3.5 GPT-4o GPT-3.5 GPT-4o
13 $0.565$ $0.581$ $0.520$ $0.537$ $0.586$ $0.592$
26 $0.576$ $0.595$ $0.515$ $0.553$ $0.595$ $0.617$
65 $0.584$ $0.622$ $0.539$ $0.584$ $0.601$ $0.636$
130 $0.631$ $0.684$ $0.579$ $0.639$ $0.647$ $0.697$
260 $0.682$ $0.689$ $0.652$ $0.657$ $0.697$ $0.701$
520 $0.687$ $0.688$ $0.654$ $0.659$ $0.700$ $0.702$
As shown in Table VIII, GPT-4o initially outperforms GPT-3.5 across all metrics (M-IoU, IoU, and $F_{2}$ Score) when the dataset size is small (e.g., with a set size of 13, GPT-4o achieves an M-IoU of 0.581 compared to 0.565 for GPT-3.5). This suggests that GPT-4o has stronger generalization capabilities with limited data, allowing it to perform better in low-resource scenarios. However, as the dataset size increases, the performance gap between GPT-4o and GPT-3.5 begins to narrow, particularly at larger set sizes (e.g., at a set size of 520, both models achieve similar M-IoU scores of 0.687 for GPT-3.5 and 0.688 for GPT-4o). This indicates that while GPT-4o excels with smaller datasets due to its ability to generalize well from fewer examples, GPT-3.5 is able to catch up as more data becomes available.
@@ VI-A Contributions
In our study, we implemented a data augmentation approach using ChatGPT-4o to fine-tune the GPT-3.5 model for identifying and highlighting key components of trainee tutors’ responses in low-resource scenarios. This approach demonstrated a significant improvement in model performance, which in turn facilitates the provision of automated feedback in tutor training programs. Notably, our results on outcome-based and effort-based praise show that even with as few as 13 human-annotated instances, the model fine-tuned with synthetically augmented data achieved performance comparable to a model fine-tuned on 65 human-annotated instances. This finding is particularly significant as it indicates a substantial reduction in the manual labor required for annotation, making the training process more efficient and scalable. These promising results suggest that the data augmentation techniques employed in this study could be leveraged to enhance model performance in other tutoring practices, such as responding to student errors and assessing students’ knowledge [29]. This potential opens up the possibility of developing automated feedback systems (such as the HAROR feedback system [61]) for different lessons in tutor training, which is crucial for enhancing the efficacy of tutor training programs and ultimately improving the overall tutoring process.
Then, we tested the generalization ability of our method by applying our data augmentation and fine-tuning method to person-based praise. The results indicate that even with a small number of labeled data, our method effectively boosts the model’s performance. The effectiveness can be attributed to the ability of our proposed method to generate high-quality synthetic data that closely mimics real-world text distributions, allowing the model to learn robust patterns despite limited labeled data. Distinct from other text data augmentation methods (e.g., synonym replacement, random word insertion), our approach leverages the semantic coherence provided by GPT-4o, ensuring that the augmented data captures the nuanced characteristics needed for accurate model training.

### v2

@@ IV-A2 Analysis of Results
The average M-IoU, IoU, $F_{2}$ score performance of the fine-tuned models, trained with the training set split with five different random seeds, is depicted in Table IV and Table V.
TABLE IV: Average Scores and Standard Errors for M-IoU, IoU, and $F_{2}$ Score (effort-based praise). 520* denotes traditional augmentation methods (e.g., synonym replacement, random word insertion) with a set size of 520.
@@
13 $0.506$ $0.024$ $0.443$ $0.011$ $0.541$ $0.027$
26 $0.501$ $0.020$ $0.440$ $0.009$ $0.547$ $0.034$
65 $0.495$ $0.021$ $0.436$ $0.009$ $0.540$ $0.025$
130 $0.551$ $0.018$ $0.488$ $0.013$ $0.615$ $0.006$
260 $0.596$ $0.008$ $0.551$ $0.004$ $0.636$ $0.012$
520 $0.601$ $0.006$ $0.554$ $0.007$ $0.652$ $0.013$
520* $0.456$ $0.036$ $0.392$ $0.014$ $0.490$ $0.027$
TABLE V: Average Scores and Standard Errors for M-IoU, IoU, and $F_{2}$ Score (outcome-based praise). 520* denotes traditional augmentation methods (e.g., synonym replacement, random word insertion) with a set size of 520.
@@
Table IV presents the scores and errors for effort-based praise across different augmented training set sizes, ranging from 13 to 520 labeled responses, while Table V details the corresponding metrics for outcome-based praise. In both cases, as the training set size increases, there is a noticeable improvement in all three metrics—M-IoU, IoU, and $F_{2}$ scores. For effort-based praise, the M-IoU score starts at 0.506 with a set size of 13 and gradually increases to 0.601 with a set size of 520, with similar trends observed for the IoU and $F_{2}$ scores, indicating that larger augmented training sets contribute to better model performance. Similarly, for outcome-based praise, the M-IoU score begins at 0.623 for the smallest set size of 13 and reaches 0.772 with the largest set size of 520, with consistent improvements also seen in the IoU and $F_{2}$ scores. These results reinforce the effectiveness of data augmentation for enhancing the model’s ability to accurately identify both effort-based and outcome-based praise. In addition to our primary methods, we employed traditional NLP data augmentation methods—random insertion, swap, deletion, and synonym replacement—to expand the original 13 responses into a dataset of 520 entries. Specifically, each original response was augmented by applying these traditional methods collectively rather than generating separate datasets for each method individually. For example, a single tutor response would undergo synonym replacement, random word insertion, swapping, and deletion in combination, producing several augmented responses simultaneously. This combined augmentation approach allowed for more extensive dataset expansion by creating more diverse variants of each original response. The augmented dataset, containing a total of 520 responses, is labeled as ’520*’ in Table IV and Table V, where we report the performance of models trained on this expanded dataset. However, these traditional data augmentation methods did not enhance performance and, in fact, resulted in worse outcomes compared to using no augmentation at all. This lack of improvement could be attributed to the introduction of incoherent or invalid sentences by these methods. For instance, the sentence “Hey Kevin, you did a great job.” was altered to “Hey, Kevin, you the did an good work,” which illustrates the potential for reduced clarity and grammatical correctness.
To statistically validate the observed improvements due to increasing the augmented dataset size, we performed the Mann-Whitney U test on key comparisons between data augmentation levels. For effort-based praise, although comparisons between the smallest sets (e.g., 13 vs. 26 and 13 vs. 65) did not show statistically significant differences (e.g., $U=13.00$, $p=0.5794$), larger increases in training data exhibited significant improvements; notably, comparing 13 vs. 260 responses yielded $U=4.00$ and $p=0.0476$, indicating significance at the 0.05 level. Similar trends were found when comparing intermediate augmentations such as 26 vs. 130 ($U=4.00$, $p=0.0476$) and 26 vs. 260 ($U=1.00$, $p=0.0079$), confirming that increasing data augmentation size significantly improves model performance.
For outcome-based praise, more pronounced statistical significance was observed. While smaller comparisons such as 13 vs. 26 ($U=9.00$, $p=0.2738$) did not reach significance, larger increments, including 13 vs. 130 ($U=2.00$, $p=0.0159$) and 13 vs. 260 ($U=0.00$, $p=0.0060$), demonstrated significant improvements. Additional comparisons such as 65 vs. 130 ($U=3.00$, $p=0.0278$) and 65 vs. 260 ($U=0.00$, $p=0.0060$) further confirmed these findings.
@@ IV-B1 Experimental Setup and Data Imbalance
As discussed in Section III-A, we have a limited dataset with only 10 labeled responses featuring person-based praise. We combine these with our existing 129 labeled responses that do not include person-based praise, which results in a total of 139 labeled responses. Due to the significant data imbalance, where the majority of the praise is effort-based or outcome-based and only a small portion is person-based, we approach this as a two-class NER problem—instead of distinguishing between effort-based praise and outcome-based praise as in RQ1, we only require the model to identify person-based praise (labeled as P) and other parts (labeled as O). If we were to fine-tune the model on all types of praise without addressing this imbalance, the model might be biased towards predicting effort-based or outcome-based praise, given their predominance in the dataset. Following the method used in RQ1, we split the dataset into a training set and a test set. The training set consists of 70 labeled responses, including 5 person-based praise responses, while the test set contains 69 labeled responses, also including 5 person-based praise responses. We then apply our data augmentation method to both the training set and the test set. The augmentation multipliers were set to $\times 1,\times 2,\times 3$, and $\times 5$. The augmentation multipliers were set to $\times 1,\times 2,\times 3$, and $\times 5$ based on the model performance. In addition, we also augmented the person-based responses in the test set by $\times 20$ to obtain more precise evaluation values.
@@ IV-B2 Generalization Performance
The results of this experiment are presented in Table VI.
TABLE VI: Average Scores and Standard Errors for M-IoU, IoU, and $F_{2}$ Score (person-based praise). Aug. Size represents Augmented Training Set Size, and Mult. represents Augmentation Multiplier.
@@
$\times 1$ 70 $0.808$ $0.014$ $0.797$ $0.016$ $0.837$ $0.009$
$\times 2$ 140 $0.897$ $0.013$ $0.895$ $0.014$ $0.924$ $0.010$
$\times 3$ 210 $0.966$ $0.017$ $0.965$ $0.013$ $0.975$ $0.006$
$\times 5$ 350 $0.969$ $0.016$ $0.969$ $0.017$ $0.975$ $0.014$
As is shown in Table VI, as the augmented training dataset size increases from 70 to 350, there is a clear upward trend across all three performance metrics: M-IoU, IoU, and $F_{2}$ Score. Specifically, the M-IoU score increases from 0.808 at a size of 70 to 0.969 at a size of 350. Similarly, the IoU score improves from 0.797 to 0.969, and the $F_{2}$ Score rises from 0.837 to 0.975. These results indicate that as we augment the training dataset, the model’s ability to identify person-based praise improves consistently across all measured metrics. This gradual improvement demonstrates the effectiveness of our data augmentation method in enhancing the model’s performance.
Specifically, when the data augmentation multiplier is set to $\times 3$ (namely when the augmented training dataset size is 210), the model achieves satisfactory performance, the M-IoU, IoU, and $F_{2}$ scores already reached 0.965 or higher, and even achieved a perfect score of 1 under certain random seeds, suggesting that the model was nearing its performance ceiling, which demonstrates the model’s robust ability to generalize to person-based praise. Beyond this multiplier, at $\times 5$, the performance plateaus, where increasing the multiplier to $\times 5$ yielded only marginal improvements, suggesting that further augmentation does not yield additional benefits.
The relatively high performance achieved with smaller augmentation multipliers (e.g., $\times 3$, namely the augmented training dataset size is 210) can be attributed to the nature of person-based praise, which often consists of straightforward adjectives describing a person’s characteristics (e.g., “You are smart.”, “You are the smartest student.”). The simplicity in the vocabulary and word types—predominantly adjectives—makes it easier for the model to recognize and tag person-based praise, in contrast to the more nuanced effort-based and outcome-based praise categories examined in RQ1. This straightforward linguistic structure allows the model to detect patterns with greater ease, as opposed to the complexity found in the other categories, where varied and context-dependent language may pose more challenges for accurate tagging.
@@ V Ablation Study
To assess whether more advanced LLMs, such as GPT-4o, provide better data augmentation performance compared to GPT-3.5, we first evaluate the performance of data augmentation using GPT-4o and GPT-3.5, keeping all other settings consistent with the previous sections. The results are presented in Table VII.
As shown in Table VII, augmenting the dataset with GPT-4o consistently outperforms augmentation with GPT-3.5 across all metrics (M-IoU and IoU Score). This improvement can be explained by GPT-4o’s ability to generate more valid and contextually appropriate synonym replacements compared to GPT-3.5. Additionally, we observed that GPT-3.5 often fails to adhere strictly to the clean answer structure we set in the prompt, leading to inconsistencies in the output format and introducing some noise into the augmented data. This issue was significantly less prevalent in outputs generated by GPT-4o, which produced more reliable and coherent augmentations. For example, when augmenting a sentence like “You did a good job,” GPT-3.5 might generate expressions such as “Sure, here is the synonym: You executed a commendable operation!” which may introduce the noise “Sure, here is the synonym:”.
TABLE VII: Average M-IoU and IoU Score of augmentation with GPT-4o and GPT-3.5. We keep the base model as GPT-3.5.
M-IoU IoU
Set Size GPT-4o GPT-3.5 GPT-4o GPT-3.5
13 $0.565$ $0.554$ $0.520$ $0.493$
26 $0.576$ $0.569$ $0.515$ $0.508$
65 $0.584$ $0.574$ $0.539$ $0.521$
130 $0.631$ $0.598$ $0.579$ $0.546$
260 $0.682$ $0.601$ $0.652$ $0.550$
520 $0.687$ $0.609$ $0.654$ $0.561$
@@ V-B Initial Training Dataset Size
In this section, we investigate the impact of different initial training dataset sizes on model performance after data augmentation. We use GPT-4o for data augmentation while keeping the base model as GPT-3.5. The results are presented in Table VIII, which shows the average M-IoU for training sets with different initial sizes when augmented to size 130, 260, and 520.
TABLE VIII: Average M-IoU for Training Sets with Different Initial Sizes after Augmentation.
@@
As shown in Table VIII, increasing the initial training set size leads to improved performance after augmentation, but the gains diminish as the augmented dataset size grows larger. For instance, starting with an initial set size of 13, the M-IoU improves from 0.631 to 0.687 as the augmented set size increases from 130 to 520. However, when starting with an initial set size of 65, the performance gains become minimal, suggesting that when the initial set size is relatively large, further augmentation may not be necessary or only a little augmentation is required. Beyond a certain point, additional augmentation yields diminishing returns. This observation indicates that while data augmentation is effective in low-resource scenarios (e.g., starting with only 13 or 26 labeled instances), its impact becomes less pronounced as more labeled data is available initially. Therefore, for small datasets, augmentation plays a crucial role in boosting model performance, but its utility decreases as the dataset grows larger.
@@ V-C Base Model
In this section, we compare the performance of GPT-3.5 and GPT-4o as base models, while keeping GPT-4o as the augmentation model. The results are shown in Table IX.
TABLE IX: Average M-IoU and IoU Score with GPT-4o and GPT-3.5 as base model. We keep augmentation model as GPT-4o.
M-IoU IoU
Set Size GPT-3.5 GPT-4o GPT-3.5 GPT-4o
13 $0.565$ $0.581$ $0.520$ $0.537$
26 $0.576$ $0.595$ $0.515$ $0.553$
65 $0.584$ $0.622$ $0.539$ $0.584$
130 $0.631$ $0.684$ $0.579$ $0.639$
260 $0.682$ $0.689$ $0.652$ $0.657$
520 $0.687$ $0.688$ $0.654$ $0.659$
As shown in Table IX, GPT-4o initially outperforms GPT-3.5 across all metrics (M-IoU, IoU, and $F_{2}$ Score) when the dataset size is small (e.g., with a set size of 13, GPT-4o achieves an M-IoU of 0.581 compared to 0.565 for GPT-3.5). This suggests that GPT-4o has stronger generalization capabilities with limited data, allowing it to perform better in low-resource scenarios. However, as the dataset size increases, the performance gap between GPT-4o and GPT-3.5 begins to narrow, particularly at larger set sizes (e.g., at a set size of 520, both models achieve similar M-IoU scores of 0.687 for GPT-3.5 and 0.688 for GPT-4o). This indicates that while GPT-4o excels with smaller datasets due to its ability to generalize well from fewer examples, GPT-3.5 is able to catch up as more data becomes available.
@@ VI-A Contributions
In our study, we implemented a data augmentation approach using ChatGPT-4o to fine-tune the GPT-3.5 model for identifying and highlighting key components of trainee tutors’ responses in low-resource scenarios. This approach demonstrated a significant improvement in model performance, which in turn facilitates the provision of automated feedback in tutor training programs. Our results show that with only 13 annotated instances, a model fine-tuned on synthetically augmented data can perform similarly to one trained on 65 annotated instances. This finding is particularly significant as it indicates a substantial reduction in the manual labor required for annotation, making the training process more efficient and scalable. These promising results suggest that the data augmentation techniques employed in this study could be leveraged to enhance model performance in other tutoring practices, such as responding to student errors and assessing students’ knowledge [28]. This suggests the potential to develop automated feedback systems for various tutor training lessons, improving program efficacy and the tutoring process overall.
Then, we tested the generalization ability of our method by applying our data augmentation and fine-tuning method to person-based praise. The results indicate that even with a small number of labeled data, our method effectively boosts the model’s performance. The effectiveness can be attributed to the ability of our proposed method to generate high-quality synthetic data that closely mimics real-world text distributions, allowing the model to learn robust patterns despite limited labeled data. Unlike other augmentation methods such as synonym replacement, our approach uses GPT-4o to maintain semantic coherence, capturing nuances crucial for effective model training.